# Federated Learning Privacy: Attacks, Defenses, Applications, and Policy Landscape — A Survey

### v1

@@ 6.2. Membership Inference Attacks
The goal of membership inference attacks is to infer whether any particular data instance has been used in the training of a specific model. If a particular data instance has been used in the training, this instance is called a member, otherwise it is a non-member. Knowing the membership of one particular data point could result in revealing private information, for example if someone’s data is known to be in a cancer dataset (used to train a cancer prediction model), then it is highly likely that this particular person has cancer.
@@
Nasr et al. (2019) presented the very first analysis on membership inference attack against neural networks under FL with access to some training members and white-box access to the trained neural networks. The adversary can be either the central server or one of the participants in the FL framework. Besides, they presented the definition of passive attacker and active attacker, where the passive attacker does the training normally and the active attacker breaks the FL protocols to improve the effectiveness of MI attacks.
For passive attackers, they showed that the attackers can take advantage of the gradients, activation maps, prediction vectors, loss and true label of one instance using local model of each user at different epochs during the training process and perform MI attacks. An attack model is trained using some known members.
On the other hand, two active attacker strategies were proposed. The first one is called gradient ascent attacker, which means that the attacker will do gradient ascent on targeted samples. By taking the action of gradient ascent, the loss of the targeted samples will be increased. For member instances, their loss will be abruptly decreased by one client, thus can be distinguished from non-members. Both central server adversary and client adversary can utilize this strategy. The second active attack method is called isolating attacker. The attacker can isolate one special target participant by creating a separated copy of the aggregated model. This is equivalent to training a model on a single machine using the data from the targeted client. To use this isolation attack, the adversary needs to be the central server.
Zari et al. (2021) proposed a computationally more efficient passive membership inference attack under FL which also requires some known members. Their attack uses as the feature vector the probabilities of the correct label under local models at different epochs, namely , where is the number of communication rounds. An attack model is trained to predict membership using known members and known non-members (which come from validation). Since the feature vector contains only one number from one epoch, this attack requires significantly less computational resources than the Nasr’s attack which uses information such as gradient and activation maps. The authors showed that using this probability feature vector can achieve higher membership inference attack accuracy than Nasr’s attack on CIFAR-100 dataset with AlexNet and DenseNet, but lower accuracy than Nasr’s attack on Purchase dataset.
@@
It is worth noting that all the membership inference attacks that do not require shadow models can be applied to FL setting as well. Shadow model is a commonly used technique in membership inferences and by this technique, the attacker is able to train models similar to the target model using data drawn from the same distribution of the training data of the target model. Thus, shadow models enable attackers to collect any information necessary for membership inference, for example, what the loss would be for one instance if this instance is not included in training.
Yeom et al. (2018) presented a theoretical analysis on privacy leakage based on prediction loss. They showed that the generalization gap between training accuracy and testing accuracy could serve as a lower bound for MI attack accuracy, as the attacker predicts member if and only if the model’s prediction is correct. Besides, they proposed a threshold attack based on prediction loss, and the attacker predicts member if and only if the prediction loss is lower than the threshold, since the loss of members is generally lower than the loss of non-members.
Inspired by Yeom et al. (2018), Song et al. (2019) proposed to adjust the prediction loss by a class-dependent value by noticing that some classes are harder to correctly classify than other classes. In addition, Song et al. (2019) proposed to use a modified prediction entropy to predict membership. The modified prediction entropy is calculated using this formula: .
@@
Hui et al. (2021) exploits the model’s prediction in a special “set” fashion. The adversary is aware of a set of members, and of non-members, and compute and , where denotes the Maximum Mean Discrepancy (MMD) (Fortet and Mourier, 1953) of two distributions of features extracted from the prediction vectors. The intuition is that if is a member, then we tend to have , and if is non-member, we tend to have . This attack only provides predicted membership, which means it is infeasible to evaluate its effectiveness at a low false positive rate.
@@ 6.3. Property Inference Attacks
Fredrikson et al. (2015) proposed the very first black-box property inference attack where the goal is to infer sensitive attributes of a given target record which is used in the training of a target model. This black box property inference attack assumes that the adversary can obtain the model’s predicted label, has knowledge of all the attributes of a targeted record (including the true label) except the sensitive attribute, has access to the marginal priors of all the attributes, and also to the confusion matrix of the target model (which is usually not known by the attacker). The adversary queries the target model by varying the sensitive attribute and obtains the predicted labels. Assuming that there is different values for the single sensitive attribute that the attacker wants to know, after querying the model multiple times with different possible values while keeping the other known attributes unchanged, the adversary computes for each possible sensitive attribute value, where (which comes from the confusion matrix) and , which is the marginal prior of i-th possible sensitive attribute value. Finally, the attack predicts the sensitive attribute value for which the value is the maximum.
@@
Mehnaz et al. (2022) proposed two new attacks: a confidence score based model inversion attribute inference attack and a label-only model inversion attack. The key intuition of the confidence score based attack is that the target model’s returned prediction is more likely to be correct and the confidence score is more likely to be higher when it is queried with a record containing the original sensitive attribute value. Thus, the value that can get correct prediction and yield lowest loss is predicted to be the correct value of the sensitive attribute. For the label-only model inversion attack, the attacker can collect a subset of target samples which satisfy the following conditions: there is only a single value of the sensitive attribute to make the prediction correct. Once this subset is successfully collected, the attacker can train an attack model to predict the value of the sensitive attribute, assuming the collected set is always correct. They also extend our attacks to the scenario where some of the other (non-sensitive) attributes of a target record are unknown to the adversary. Moreover, they empirically demonstrate the disparate vulnerability of model inversion attacks, i.e., specific groups in the training dataset (grouped by gender, race, etc.) could be more vulnerable to model inversion attacks. In (Dibbo et al., 2023), Dibbo et al. refined the method above by proposing a new strategy to collect training data for the attack model. Instead of using real world data, they proposed to use generated random instances and change the value of the sensitive attribute to query the target model. Moreover, for each random instance, it is only accepted when each different sensitive attribute results in a different predicted label. Once this random instance set is gathered, an attack model is trained to predict the sensitive attribute for instances that the attacker is interested in.
@@ 6.4. Model Extraction Attacks
In model extraction attacks (Oliynyk et al., 2023; Gong et al., 2020), an attacker can establish a substitute model with almost the same functionalities as the target victim model via simply querying the victim model. While these attacks are not a direct threat to client data privacy in FL, they pose a threat against the server, stealing either the parameters, hyper parameters, architecture, or functionality of the model (Tramèr et al., 2016; Wang and Gong, 2018; Orekondy et al., 2019). As such, model extraction attacks serve as a privacy threat for the intellectual property (i.e., the trained model) of a company. This is a practical impediment to the widespread deployment of FL as the trained model at the server is considered a precious asset. Hence, threats that allow an adversary to steal that asset have to be defended with a high degree of certainty.
Further, model extraction attacks can serve as a springboard to attacks that violate client data privacy. For example, the adversary can use the proxy model created using model extraction to launch attacks that translate well to the original model. Due to the transferability property of such attacks (the attacks against the proxy model are accurate against the original model) (Wang and He, 2021), the adversary can do a membership inference attack to determine which client data items were used in creating the model. This is worrisome as it has been shown recently that the best previous defense of perturbing the results of the query by the adversary as it is trying to construct the proxy model is still vulnerable to model extraction (Chen et al., 2023a). The basic idea is that this attack module, which can be integrated with any model extraction attack, can determine the difference in distribution between the unperturbed and perturbed query results and ”reset” that difference.
@@ 7. Privacy Defenses
| Defense class | Summary | Citations | Effective against | Resources | ||||||||||||||||||||||||||||
(Dwork, 2008; Abadi et al., 2016; Dong et al., 2019; Koskela et al., 2020; Steinke et al., 2023) |
(Bonawitz et al., 2017; Bell et al., 2020; Kadhe et al., 2020; Zhao and Sun, 2022; So et al., 2022) |
(Cheon et al., 2017; Zhang et al., 2020; Fang and Qian, 2021) |
@@ 7.1. Differential Privacy
DP based methods can provide a theoretical upper-bound on the effectiveness of any MI attack against any instance. Unfortunately, achieving a meaningful theoretical guarantee (e.g., with a resulting ) requires the usage of very large noises and thus resulting a huge utility loss. However, model trainer could use much smaller noises in DP-SGD. While doing this fails to provide a meaningful theoretical guarantee (the value would be too large), this can nonetheless provide empirical defense against MI attacks to protect privacy.
@@ 7.2. Secure Aggregation
Secure aggregation (SA) (Bonawitz et al., 2017) enables the server to aggregate local model updates from a number of users, without observing any of their model updates in the clear. Specifically, SA protocols ensure that (1) the server and any set of users do not learn any information about the local dataset of any user from the encrypted model updates in the information theoretic sense; (2) the server only learns the aggregated model; (3) correct decoding of the aggregated model in the presence of users dropout.
The state-of-the-art for secure aggregation protocols in FL is to use additive masking to protect the privacy of individual models (Bonawitz et al., 2017; Bell et al., 2020; So et al., 2021a; Kadhe et al., 2020; Zhao and Sun, 2022; So et al., 2022, 2021b). SecAgg (Bonawitz et al., 2017) is the first practical protocol proposed by Google for FL that is resilient for both user failures or dropouts and collusion between users as well as the server. Particularly, SecAgg as depicted in Figure 7 leverages pairwise random seed , for , to be generated between each pair of users for masking the model updates. Each pair of users generate the pairwise random seed by using key agreement protocol (e.g., Diffie-Hellman (Diffie and Hellman, 1976)), such that the random seed is a function of the public key of user and the private key of user . In addition, user creates a private random seed to prevent the privacy breaches that may occur if user is only delayed rather than dropped, in which case the pairwise masks alone are not sufficient for privacy protection. For handling users failures or dropouts, the private key and the private random seed of each user are secret shared among all other users, and can be reconstructed by the server if any user drops during the protocol, which allows for a more resilient recovery protocol against user dropouts. User then masks its model as , where PRG is a pseudo random generator, and sends it to the server. Finally, to remove the masks that involves the dropped users, the server asks the set of survived users for the shares of the private keys of the dropped users and the shares of the private seeds of the survived users. Figure 7 gives an example for applying SA in the FL setting of three users. The cost of constructing and sharing the masks in SecAgg, however, scales with respect to with corresponding to the number of users, which takes the majority of execution time.
@@
Recently, (So et al., 2022) reduced the overhead of model aggregation in SecAgg by proposing LightSecAgg that relies on using private Maximum Distance Separable (MDS) codes (Roth and Lempel, 1989; Yu et al., 2019; Tang et al., 2021). The reduction in complexity achieved by LightSecAgg is based on using one-shot aggregate-mask reconstruction of the surviving users instead of pairwise random-seed reconstruction of the dropped users that reduces the reconstruction complexity to compared to and for SecAgg+ and SecAgg, respectively. LightSecAgg is composed of three main phases as shown in Figure 8. First, each user partitions its local random mask to pieces and creates encoded masks via private MDS code to provide robustness against dropped users and privacy against colluding users. Each user sends one of the encoded masks to one of the other users for the purpose of one-shot recovery. Second, each user uploads its masked local model to the server. Third, the server reconstructs the aggregated masks of the surviving users to recover their aggregate of models. Each surviving user sends the aggregated encoded masks to the server. After receiving enough aggregated encoded masks from the surviving users, the server recovers the aggregate-mask and the desired aggregate-model.
Another interesting research direction of SA in FL has been to further analyze its formal privacy guarantees either over multi-rounds (So et al., 2023) or from the aggregated model (Elkordy et al., 2022). In particular, while secure aggregation protocols have provable privacy guarantees at any single round, in the sense that no information is leaked beyond the aggregate model at each round, the privacy guarantees do not extend to attacks that span multiple training rounds. The authors in (So et al., 2023) have shown that the individual model may be reconstructed by leveraging the participation information and the aggregate models across multiple rounds. Assuming a scenario where the local updates do not change significantly over time (e.g., models start to converge), i.e., for all and . Then the server can single out a model, say , even if a secure aggregation protocol is employed at each round. Instead of random users participation, the authors managed to resolve this limitation of secure aggregation by developing a structured user selection strategy that guarantees long-term privacy while taking into account the fairness in user selection and average number of participating users, and showed that provides a trade-off between long-term privacy and the convergence rate.
SA design ensures that the server only learns the aggregated model on each round, Figure 9. However, even with these SA guarantees on individual updates, it is not yet fully understood how much privacy is guaranteed in FL using SA, since the aggregated model update from all users may still leak information about an individual user’s local dataset (Pasquini et al., 2022; Zhao et al., 2023b). The work in (Elkordy et al., 2022) measures the amount information the aggregated model leak about the local dataset of an individual user. They provide theoretical bounds on the mutual information (MI) privacy leakage in theory and demonstrate through an empirical study that these bounds hold in practice. Their key observations indicate that when employing FL in conjunction with SA, several significant trends emerge. First, they found that the MI privacy leakage decreases at a rate of , where is the number of users participating in FL with SA. Second, they noted that increasing model size does not lead to a linear increase of MI privacy leakage, and the MI privacy leakage only linearly increases with the rank of the covariance matrix of the individual model update. Lastly, using larger batch size during local training can help to reduce the MI privacy leakage. As highlighted, the privacy metric used in that work (mutual information) measures the on-average privacy leakage, without providing any privacy guarantees for worse-case scenarios. Measuring how much differential privacy SA can achieve might be an interesting direction to be explored.

### v2

@@ 6.2. Membership Inference Attacks
The goal of membership inference attacks is to infer whether any particular data instance has been used in the training of a specific model. If a particular data instance has been used in the training, this instance is called a member, otherwise it is a non-member. Knowing the membership of one particular data point could result in revealing private information, for example, if someone’s data is known to be in a cancer dataset (used to train a cancer prediction model), then it is highly likely that this particular person has cancer. Figure 4 summarizes the process of membership inference attacks.
@@
Nasr et al. (2019) presented the very first analysis of the membership inference attack against neural networks under FL with access to some training members and white-box access to trained neural networks. The adversary can be either the central server or one of the participants in the FL framework. They also presented the definition of passive attacker and active attacker, where the passive attacker does the training normally and the active attacker breaks the FL protocols to improve the effectiveness of MI attacks.
For passive attackers, they showed that attackers can take advantage of the gradients, activation maps, prediction vectors, loss, and true label of a single instance using local model of each user at different epochs during the training process and perform MI attacks. An attack model is trained using some known members.
On the other hand, two active attacker strategies were proposed. The first is the gradient ascent attacker, which means that the attacker will apply gradient ascent on targeted samples. By taking the action of gradient ascent, the loss of the targeted samples will increase. For member instances, their loss will be abruptly decreased by one client, thus allowing distinguishability from non-members. Both a central server adversary and a client adversary can utilize this strategy. The second active attack method is called the isolating attacker. The attacker can isolate one special target participant by creating a separate copy of the aggregated model. This is equivalent to training a model on a single machine using the data from the targeted client. To use this isolation attack, the adversary would be a malicious central server.
Zari et al. (2021) proposed a computationally more efficient passive membership inference attack under FL which also requires some known members. Their attack uses as the feature vector the probabilities of the correct label under local models at different epochs, namely , where is the number of communication rounds. An attack model is trained to predict membership using known members and known non-members (which come from validation). Since the feature vector contains only one number from one epoch, this attack requires significantly less computational resources than the Nasr’s attack which uses information such as gradient and activation maps. The authors showed that using this probability feature vector can achieve higher membership inference attack accuracy than Nasr’s attack on CIFAR-100 dataset with AlexNet and DenseNet, but lower accuracy on Purchase dataset.
@@
It is worth noting that all the membership inference attacks that do not require shadow models can be applied to FL setting as well. Shadow models are a commonly used technique in membership inference attacks, and allow the attacker =to train models similar to the target model using data drawn from the same distribution of the training data of the target model. Thus, shadow models enable attackers to collect any information necessary for membership inference, for example, what the loss would be for one instance if this instance is not included in training.
Yeom et al. (2018) presented a theoretical analysis on privacy leakage based on prediction loss. They showed that the generalization gap between training accuracy and testing accuracy could serve as a lower bound for MI attack accuracy, as the attacker predicts the member if and only if the model’s prediction is correct. Besides, they proposed a threshold attack based on prediction loss, and the attacker predicts member if and only if the prediction loss is lower than the threshold, since the loss of members is generally lower than the loss of non-members.
Inspired by Yeom et al. (2018), Song et al. (2019) proposed to adjust the prediction loss by a class-dependent value, noticing that some classes are harder to correctly classify than other classes. In addition, Song et al. (2019) proposed using a modified prediction entropy to predict membership. The modified prediction entropy is calculated using this formula: .
@@
Hui et al. (2021) exploits the model’s prediction in a special “set” fashion. The adversary is aware of a set of members, and of non-members, and compute and , where denotes the Maximum Mean Discrepancy (MMD) (Fortet and Mourier, 1953) of two distributions of features extracted from the prediction vectors. The intuition is that if is a member, then we tend to have , and if is non-member, we tend to have . This attack only provides predicted membership, which means that it is infeasible to evaluate its effectiveness at a low false positive rate.
@@ 6.3. Property Inference Attacks
Fredrikson et al. (2015) proposed the very first black-box property inference attack where the goal is to infer sensitive attributes of a given target record which is used in the training of a target model. This black box property inference attack assumes that the adversary can obtain the model’s predicted label, has knowledge of the attributes of a targeted record (including the true label) except the sensitive attribute, has access to the marginal priors of the attributes, and also to the confusion matrix of the target model (which is usually not known by the attacker). The adversary queries the target model by varying the sensitive attribute and obtains the predicted labels. Assuming that there is different values for the single sensitive attribute that the attacker wants to know, after querying the model multiple times with different possible values while keeping the other known attributes unchanged, the adversary computes for each possible sensitive attribute value, where (which comes from the confusion matrix) and , which is the marginal prior of i-th possible sensitive attribute value. Finally, the attack predicts the sensitive attribute value for which the value is the maximum.
@@
Mehnaz et al. (2022) proposed two new attacks: a confidence score based model inversion attribute inference attack and a label-only model inversion attack. The key intuition of the confidence score based attack is that the target model’s returned prediction is more likely to be correct and the confidence score is more likely to be higher when it is queried with a record containing the original sensitive attribute value. Thus, the value that can get correct prediction and yield lowest loss is predicted to be the correct value of the sensitive attribute. For the label-only model inversion attack, the attacker can collect a subset of target samples which satisfy the following conditions: there is only a single value of the sensitive attribute to make the prediction correct. Once this subset is successfully collected, the attacker can train an attack model to predict the value of the sensitive attribute, assuming that the collected set is always correct. They also extend the attacks to the scenario where some of the other (non-sensitive) attributes of a target record are unknown to the adversary. Moreover, they empirically demonstrate the disparate vulnerability of model inversion attacks, i.e., specific groups in the training dataset (grouped by gender, race, etc.) could be more vulnerable to model inversion attacks. Dibbo et al. (2023) refined the above method by proposing a new strategy to collect training data for the attack model. Instead of using real-world data, they proposed using random generated instances and changing the value of the sensitive attribute to query the target model. Moreover, for each random instance, it is only accepted when each different sensitive attribute results in a different predicted label. Once this random instance set is gathered, an attack model is trained to predict the sensitive attribute for instances that the attacker is interested in.
@@ 6.4. Model Extraction Attacks
In model extraction attacks (Oliynyk et al., 2023; Gong et al., 2020), an attacker can establish a substitute model with almost the same functionalities as the target victim model by querying the victim model. Although these attacks are not a direct threat to client data privacy in FL, they pose a threat against the server by stealing the parameters, hyper parameters, architecture, or functionality of the model (Tramèr et al., 2016; Wang and Gong, 2018; Orekondy et al., 2019). As such, model extraction attacks serve as a privacy threat for the intellectual property (i.e., the trained model) of a company. This is a practical impediment to the widespread deployment of FL as the trained model of the server is considered a valuable asset. Hence, threats that allow an adversary to steal that asset should be defended with a high degree of certainty.
Model extraction attacks can also serve as a springboard for attacks that violate client privacy. For example, the adversary can use the proxy model created using model extraction to launch attacks that translate well to the original model. Due to the transferability property of such attacks (the attacks against the proxy model are accurate against the original model) (Wang and He, 2021), the adversary can use a membership inference attack to determine which client data items were used in creating the model. This is worrisome, as it has been recently shown that the best previous defense of perturbing the results of the query by the adversary as it is trying to construct the proxy model is still vulnerable to model extraction (Chen et al., 2023a). The basic idea is that this attack module, which can be integrated with any model extraction attack, can determine the difference in distribution between the unperturbed and perturbed query results and ”reset” that difference.
@@ 7. Privacy Defenses
This section discusses the core privacy defenses in FL, which are summarized in Table 4 below.
| Defense class | Summary | Citations | Effective against | Resources |
| Differential privacy | Prevents privacy leakage of any individual training sample by adding noise or otherwise perturbing model updates. | (Dwork, 2008; Abadi et al., 2016; Dong et al., 2019; Koskela et al., 2020; Steinke et al., 2023) | Data reconstruction, membership inference (suitable for cross-silo and cross-device) | Clipping and noise added based on a privacy budget. Noise injection can be done at each client or at the aggregated-update level. |
| Secure aggregation | Individual updates are encrypted so the server only sees an aggregate update, never a single client’s unencrypted data. | (Bonawitz et al., 2017; Bell et al., 2020; Kadhe et al., 2020; Zhao and Sun, 2022; So et al., 2022) | Data reconstruction (suitable mainly for cross-device) | Requires additional client-to-client communication for encryption masks. Typically lower overhead than homomorphic encryption. |
| Homomorphic encryption | Clients send encrypted updates that the server can partially operate on without decrypting. The server then returns an encrypted aggregate to clients, never seeing raw updates. | (Cheon et al., 2017; Zhang et al., 2020; Fang and Qian, 2021) | Data reconstruction, membership inference, property inference (suitable for cross-silo) | High communication and computation overhead for encryption/decryption. Generally more practical for cross-silo FL. |
@@ 7.1. Differential Privacy
Limitations. DP based methods can provide a theoretical upper-bound on the effectiveness of any MI attack against any instance. However, achieving a meaningful theoretical guarantee (e.g., with a resulting ) requires the usage of very large noises and thus results in a significant drop in accuracy. (Abadi et al., 2016) gives an example for CIFAR-10 using a two convolutional layer, two linear layer network. Without differential privacy, the network can achieve 86% test accuracy. However, when are applied, the testing accuracy drops to 67%, 70%, and 73% respectively. Larger values (less noise) result in a smaller accuracy drop and have still been shown to provide some empirical defense against MI attacks (Li et al., 2021). However, we once again highlight that such large values give up on the meaningful theoretical guarantees of DP-SGD.
DP-SGD also incurs additional time and space overhead. For example, in order to perform gradient clipping, one has to store per-sample gradients. Recent efforts to improve the privacy-utility tradeoff in DP-SGD, e.g., (De et al., 2022; Sander et al., 2023), introduce additional computational overhead.
@@ 7.2. Secure Aggregation
Secure aggregation (SA) (Bonawitz et al., 2017) enables the server to aggregate local model updates from a number of users, without observing any of their model updates in the clear. Specifically, SA protocols ensure that (1) the server and any set of users do not learn any information about the local dataset of any user from the encrypted model updates in the information theoretic sense; (2) the server only learns the aggregated model; (3) correct decoding of the aggregated model in the presence of user dropout.
The state-of-the-art for secure aggregation protocols in FL is to use additive masking to protect the privacy of individual models (Bonawitz et al., 2017; Bell et al., 2020; So et al., 2021a; Kadhe et al., 2020; Zhao and Sun, 2022; So et al., 2022, 2021b). SecAgg (Bonawitz et al., 2017) is the first practical protocol proposed by Google for FL that is resilient for both user failures or dropouts and collusion between users as well as the server. Particularly, SecAgg, as depicted in Figure 6, leverages pairwise random seed , for , to be generated between each pair of users for masking the model updates. Each pair of users generate the pairwise random seed by using a key agreement protocol (e.g., Diffie-Hellman (Diffie and Hellman, 1976)), such that the random seed is a function of the public key of user and the private key of user . In addition, user creates a private random seed to prevent the privacy breaches that may occur if user is only delayed rather than dropped, in which case the pairwise masks alone are not sufficient for privacy protection. For handling users failures or dropouts, the private key and the private random seed of each user are secret shared among all other users, and can be reconstructed by the server if any user drops during the protocol, which allows for a more resilient recovery protocol against user dropouts. User then masks its model as , where PRG is a pseudo random generator, and sends it to the server. Finally, to remove the masks that involves the dropped users, the server asks the set of survived users for the shares of the private keys of the dropped users and the shares of the private seeds of the survived users. Figure 6 gives an example for applying SA in the FL setting of three users. The cost of constructing and sharing the masks, scales with respect to with corresponding to the number of users, which takes the majority of execution time.
Scalability. As the number of clients increases, defenses such as differential privacy or homomorphic encryption typically do not have a corresponding increase in the computational or communication overheads added (they are relatively fixed regardless of the number of clients). However, SA in particular has an increase in the communication costs as the number of clients in FL increase. This largely comes from the necessity of clients to communicate among themselves to ensure that the encryption works as intended (aggregation of SA encryption masks will cancel out).
@@
Additional directions. Another interesting research direction of SA in FL has been to further analyze its formal privacy guarantees either over multi-rounds (So et al., 2023) or from the aggregated model (Elkordy et al., 2022). In particular, while secure aggregation protocols have provable privacy guarantees at any single round, in the sense that no information is leaked beyond the aggregate model at each round, the privacy guarantees do not extend to attacks that span multiple training rounds. The authors in (So et al., 2023) have shown that the individual model may be reconstructed by leveraging the participation information and the aggregate models across multiple rounds.